# Overcoming the Limitations of Layer Synchronization in Spiking Neural Networks

### v1

@@ 3.1.2 Vectorized network asynchrony
The event-driven (neuron state) update rules for network asynchrony as introduced in the previous section can be vectorized by selecting multiple spikes for processing at the same time. This allows us to consider the entire spectrum of possibilities between full layer synchronization at one extreme, and “complete” asynchrony at the other extreme, where each spike event is processed entirely independently of all others. Additionally, vectorization makes acceleration possible by exploiting the parallelization features and vector pipelines of accelerators, where these models execute, leading to pragmatic simulation of network asynchrony.
During the simulation, the states of all the $N=\sum_{l=1}^{L}N^{(l)}$ neurons in the network are stored in vectors. The vector $\mathbf{x}\in\mathbb{R}^{N}$ tracks the computed input currents for the neurons, $\mathbf{u}\in\mathbb{R}^{N}$ the membrane potential of the neurons, $\mathbf{s}\in\mathbb{N}^{N}$ the emitted spikes queue, and $\mathbf{c}\in\mathbb{N}^{N}$ which neurons have spiked in the current forward pass. For any of those vectors, the indices from $\sum_{k=1}^{l-1}N^{(k)}$ to $\sum_{k=1}^{l}N^{(k)}$ represent the values for the neurons in layer $l$, for $1\leq l\leq L$ where $N^{(0)}=0$.
Algorithm 1 outlines the processing during a forward pass (propagation of the spikes of an input frame through the network). Before starting, the input spikes from time $t_{0}$ to $t_{1}$ are grouped, resulting in an input vector $\mathbf{s_{\text{in}}}\in\mathbb{N}^{N_{\text{in}}}$ where $N_{\text{in}}$ is the number of input features. Then, the input spikes are passed as one single batch, triggering a forward pass. Each forward pass consists of forward steps (the code within the while loop), which update the state of the neurons based on the spikes selected for propagation. A complete list of parameters can be found in section A.5.
@@
$\mathbf{s}\leftarrow\mathbf{0}^{N}$triangleright Vector with zeros of length $N$
@@
while ($\text{Sum}(\mathbf{x})\neq 0$ or $\text{Sum}(\mathbf{s})\neq 0$) and $\neg$¬EarlyStop($\mathbf{s}$do
@@
     $\mathbf{s}\leftarrow\mathbf{s}+\mathbf{s_{\text{new}}}$triangleright Enqueue new spikes
     $\mathbf{c}\leftarrow\mathbf{c}+\mathbf{s_{\text{new}}}$triangleright Update spike count
@@
     $\mathbf{s}\leftarrow\mathbf{s}-\mathbf{s_{\text{selected}}}$triangleright Dequeue selected spikes
@@
The forward pass always ends when all spike activitations have been processed, a stop condition referred to as "On spiking done". Additionally, EarlyStop is used to interrupt a forward pass if any neuron in the output layer has spiked zero, one, or more (default is one) forward steps ago. This is termed the "On output" stop condition.
The SelectSpikes function defines how to select a subset of the emitted spikes for propagation. The function selects $F$ spikes at a time, or less if there are less than $F$ remaining spikes ready to be propagated. The method of spike selection is determined by a scheduling policy. For the experiments, two policies are considered. The first, Random Scheduling (RS), randomly picks spikes from the entire network. The second, Momentum Scheduling (MS), prioritizes spikes from neurons based on their membrane potential upon exceeding their threshold.
@@
The network architecture is defined by the InputLayerForward and NetworkLayersForward functions. These functions compute the values of synaptic currents from spikes. For a network with only fully-connected layers, this follows from equation 3.
@@
Neuromorphic processors come in various architectures, but most of them have a template architecture that involves interconnecting many tiny processing cores with each other. This design enables a scalable architecture that supports fully distributed memory and compute systems. In this template, each processing core has its own small synchronous domain, but the cores are asynchronous among each other. For instance, the architectures of Loihi [10], SENECA [62], SpiNNaker [39], Speck [6], TrueNorth [1], POETS [50] and $\upmu$Brian [53] all follow this template. Our simulator imitates the behavior of a generalized neuromorphic processor by simulating the asynchronous interactions of several tiny synchronous domains and can be further customized to reproduce more intrinsic behavior of many of those processors (see table 4).
@@ 3.2 Training asynchronous SNNs
We use backpropagation to estimate the gradients of a loss function with respect to the weights. If multiple forward passes are needed, such as for sequential or temporal data, then Backpropagation Through Time (BPTT) can be used for training. Given that spike trains are temporal data as well, BPTT also applies to SNNs, and by extension, to the research presented here.
To address the issue of the non-differentiability of the threshold function, the surrogate gradient method is used [64]. An arctan function [13], (see section A.3 for details) provides a continuous and smooth approximation of the threshold function.
Class prediction is based on softmaxed spike counts over time, as described in section A.4. The loss is minimized using the Adam optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
@@ 3.2.1 Unlayered backpropagation
Training of SNNs with per-layer synchronization is by now straightforward using backpropagation [9] so long as the forward processing is differentiable. We refer to this as "layered backpropagation".
The vectorized network asynchronous processing approach is differentiable as well, and can be used with backpropagation. We refer to this as "unlayered backpropagation". Combined with BPTT unrolling, this method implies a two level unrolling. In the outer level, unrolling is based on discretization of time in timesteps (when new inputs are provided); within the inner level, unrolling is based on the $F$-grouping of spikes, given a scheduling policy, and the activity of neurons up-until the output is read. For a single backward pass, which can be used in BPTT in a similar way as the layered backpropagation equivalent, it is given that:
@@
where $t$ is the time(step) of the forward pass, $W$ can be a weight on any of the synapses, $L_{t}$ is the loss, $c_{t}$ is the spike count at the end of the forward pass, and $s_{i}$ are the spikes in the queue at the end of the forward step $i$, with $N_{t}$ forward steps in total within the forward pass (referring to concept of (micro)-time in 3.1.1). The number of steps scales linearly with the number of spikes processed in the forward pass. It is important to understand that due to asynchronous processing the number of spikes processed until the evaluation of the loss function may be (well) less than the total number of spikes emitted. Since for every forward step, the computations are repeated, the time complexity scales linearly with the number of forward steps: $O(N_{t})$. The same applies to the space complexity.
During the backward pass, the spikes which are not selected can skip the computation $f$ for that step:
@@
Skip connections have been researched in deep ANNs and identified as a major contributor to the stability of the training process [38]. This may apply to the skip in unlayered backpropagation as well. To what extent this is the case is not explored in this work.
@@ 4.1 Experimental setup
We experimented with trained SNN models for asynchronous execution in three common benchmarking datasets: N-MNIST [37], SHD [8], and DVS gesture [2]. Each of them has different structure. N-MNIST has purely spatial structure, SHD purely temporal, and DVS gesture combines both spatial and temporal (i.e. input framing in DVS gesture is done such that an entire gesture motion and contour is not revealed within a single frame). More details on the datasets, can be found in section A.6.1. Table 1, summarizes the parameterization of the experiments and the reported results. The network architecture and hyperparameters are given in table 5. State-of-the art performance for these tasks can be achieved with reasonably shallow and wide models. We chose however to train narrow, but deeper network architectures so that the effects of absence of layer synchronization can be revealed in the comparison (our aim is not to compete the state-of-the art). In section A.7, details on how the forward group size, refractory dropout, and momentum noise during training affect accuracy are provided.
@@
Stop condition Forward pass terminates: either when all network activity drains ("On spiking done") or one forward step after the first spike is emitted by the output layer ("On output"). The default is "On spiking done" during training and "On output" during inference.
@@ 4.2 Network asynchrony increases neuron reactivity
Interestingly while one expects that more reactive neurons imply higher activation density, surprisingly this is not the case for the models trained for asynchronous inference! In figure 2 (bottom row) we see that if we wait for the network to “drain” of spike activity the total number of spikes will indeed be higher, but if the forward pass terminates as soon as a decision is made, asynchronous models are consistently sparser. This is because asynchronous processing allows spike activity to freely flow through to the output and not be blocked at every layer for synchronization.
4.3 Unlayered backpropagation increases accuracy and sparsity
Network asynchrony negatively affects the performance of the models trained with layered backpropagation in all three datasets (table 2). This is likely in general the “Achilles’ heel” of neuromorphic AI today. However, we observe that the accuracy loss is remediated when using unlayered backpropagation, which also significantly increases sparsity (by about $2$x). In fact, the accuracy of models trained and executed asynchronously is consistently superior under both scheduling policies (with momentum scheduling having consistently an edge). This result alone suffices to attest that neuromorphic AI is competitive and more energy efficient.
@@ 4.3 Unlayered backpropagation increases accuracy and sparsity
Figure 3: Accuracy as function of forward steps after the first spike in the output layer. Models trained with layer synchronization or random scheduling, are processed with random scheduling. Models trained with momentum scheduling, are processed with momentum scheduling. Given that $F=8$, each extra forward step processes another 8 spikes, assuming enough spikes are available. Dashed lines show the accuracy after all spike activity has been “drained” out of the network.
4.4 Unlayered backpropagation decreases latency
The final and equally important result that we report is that under asynchronous network processing, models trained with unlayered backpropagation have significantly lower latency than those trained with layer synchronization. Assuming that, without parallel processing (worst case), latency is a function of the number of spikes processed until a decision is made, figure 4 shows a distribution of the inference latencies across the entire test-set. It should be clear by now that this is because “the important spikes” in these models quickly reach the output layer, uninhibited by layer synchronization.
@@ 4.4 Unlayered backpropagation decreases latency
4.5 Unlayered backpropagation is resource-intensive
The current implementation of unlayered backpropagation is costly when training with a smaller $F$ and a larger dataset. When $F$ is halved, the time and memory requirements approximately double, consistent with the linear time and space complexity discussed in section 3.2.1. The detailed results can be found in section A.8.
5 Discussion
We have introduced a variation of backpropagation training that does not assume per layer synchronization, and we use it to train models. Under asynchronous processing, these models exceed the performance of conventionally trained models and deliver the anticipated energy and latency efficiency from event-driven execution. Our findings show that unless we revise our training methods (taking into account asynchronous processing dynamics), and designs choices of neuromorphic accelerators (layer synchronization primitives considered harmful), SNNs and brain-inspired computing will likely fail to deliver either in performance or efficiency.
This study has merely scratched the surface. A key realization (that confirms neurobiology) is that trainable dynamics related to network asynchrony allow neurons (and models overall) to use partial input for decision making. The extent and generality of this effect on network architecture/structure, neuron models, and data types deserves further exploration. Our results also encourage a dynamics driven design space exploration for neuromorphic processors with regard to event processing scheduling and granularity of synchronization primitives (or lack thereof), which is almost absent today. The third exploration dimension is undoubtedly the resource efficiency and scalability of unlayered backpropagation (or analogous training framework) to facilitate very big models and datasets.

### v2

@@ 3.1.2 Vectorized network asynchrony
The event-driven (neuron state) update rules for network asynchrony as introduced in the previous section can be vectorized by selecting a number of spikes for processing them simultaneously. This allows us to consider the entire spectrum of possibilities between per-layer synchronization at one extreme (by assuming a vector size equal to a layer size), and “complete” asynchrony at the other extreme, where each spike event is processed entirely independently of all others. Additionally, vectorization makes acceleration possible by exploiting the parallelization features and vector pipelines of accelerators, where these models execute, leading to pragmatic simulation of network asynchrony.
During the simulation, the states of all the $N=\sum_{l=1}^{L}N^{(l)}$ neurons in the network are stored in vectors. Vector $\mathbf{x}\in\mathbb{R}^{N}$ tracks the computed input currents for the neurons, $\mathbf{u}\in\mathbb{R}^{N}$ the membrane potential of the neurons, $\mathbf{s}\in\mathbb{N}^{N}$ the emitted spikes awaiting processing, and $\mathbf{c}\in\mathbb{N}^{N}$ which neurons have spiked in the current forward pass. For any of those vectors, the indices from $\sum_{k=1}^{l-1}N^{(k)}$ to $\sum_{k=1}^{l}N^{(k)}$ represent the values for the neurons in layer $l$, for $1\leq l\leq L$ where $N^{(0)}=0$.
Algorithm 1 outlines the processing during a forward pass (propagation of the spikes of an input frame through the network). Input spikes are available in $\mathbf{s_{\text{in}}}\in\mathbb{N}^{N_{\text{in}}}$ (not to be confused with $\mathbf{s}$) where $N_{\text{in}}$ is the number of input features. Each forward pass consists of forward steps (the code within the while loop), which update the state of a set of neurons based on the spikes selected for propagation. A complete list of parameters can be found in section A.5.
@@
$\mathbf{s}\leftarrow\mathbf{0}^{N}$ $\triangleright$ Vector with zeros of length NN
@@
[…]
@@
  $\mathbf{s}\leftarrow\mathbf{s}+\mathbf{s_{\text{new}}}$ $\triangleright$ Enqueue new spikes
  $\mathbf{c}\leftarrow\mathbf{c}+\mathbf{s_{\text{new}}}$ $\triangleright$ Update spike count
@@
  $\mathbf{s}\leftarrow\mathbf{s}-\mathbf{s_{\text{selected}}}$ $\triangleright$ Dequeue selected spikes
@@
The forward pass ends either when all spike activations have been processed ("On spiking done" stop condition), or when any neuron in the output layer has spiked one or more (default is one) forward steps ago ("On output" stop condition). The latter is checked in the EarlyStop function.
The SelectSpikes function defines how to select a subset of the emitted spikes for propagation. The function selects $F$ spikes at a time, or less if there are less than $F$ remaining spikes ready to be propagated. The method of spike selection is determined by a scheduling policy.
For the experiments here two policies are exercised. The intent is that different scheduling strategies can be tested for their effectiveness in capturing tractably asynchronous processing dynamics. The first, Random Scheduling (RS), randomly picks spikes from the entire network. The second, Momentum Scheduling (MS), prioritizes spikes from neurons based on their membrane potential upon exceeding their threshold.
@@
The network architecture is defined by the InputLayerForward and NetworkLayersForward functions. These functions compute the values of synaptic currents from spikes. This follows from equation 3.
@@
Neuromorphic processors come in a number of variations, but most of them have a template architecture that interconnects many tiny processing cores with each other. This design enables a scalable architecture that supports fully distributed memory and compute systems. In this template, each processing core has its own small synchronous execution domain, but the cores operate asynchronously among each other. Architectures such as those cited in the Section 2 all follow this template. Our asynchronous processing simulation environment captures the intrinsics of the synchronous domain (e.g. in the forward group size $F$) while simulating the asynchronous interactions among them (e.g. scheduling policy) and can be also configured to reproduce more specialized intrinsic behaviors of many of those processors (table 5).
@@ 3.2 Training asynchronous SNNs
We use backpropagation to train the model weights, and specifically when stimulous is presented across multiple timesteps (forward passes), such as for sequential or temporal data, then this is Backpropagation Through Time (BPTT). Following common practice to address the non-differentiability of the threshold function, the surrogate gradient method is used Zenke & Vogels (2021). Specifically the arctan function Fang et al. (2021), (see section A.3 for details) provides a continuous and smooth approximation of the threshold function.
Class prediction is based on the softmaxed membrane potentials (for CIFAR-10) or spike counts over time (for the other datasets) of the neurons in the output layer, as described in section A.4. The loss is minimized using the Adam optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
@@ 3.2.1 Unlayered backpropagation
We refer to “conventional” SNN training with per-layer synchronization using backpropagation Dampfhoffer et al. (2023), as "layered backpropagation".
The vectorized network asynchronous processing approach is differentiable as well, and can be used with backpropagation. We refer to this as "unlayered backpropagation". Combined with BPTT unrolling, this method implies a two level unrolling. At the outer level, unrolling is based on discretization of time and thus the input across (as usually fixed number of) timesteps. At the inner level, unrolling is based on the $F$-grouping (and vectorized processing) of spikes in forward steps, subject to the scheduling policy, applied to the emitted spikes by neurons; from the beginning of the current timestep until the output is read.
Note, that the dependence on the scheduling policy, applied on a variable number of activations (across timesteps and data samples), in $F$-sized groups is the fundamental difference from layered backpropagation, and leads to different gradient state being built up in the computation graph. This state now captures the dynamics of asynchronous processing. For a single backward pass in a timestep of BPTT, which is applied in a similar way as the layered backpropagation equivalent, it is given that:
@@
where $t$ is the time(step) of the forward pass, $W$ refers to the trainable weights, $L_{t}$ is the loss, $c_{t}$ is the spike count at the end of the forward pass, and $s_{i}$ is the emitted spikes vector at the end of the forward step $i$. The overall gradient (state) depends on the total number of forward steps $N_{t}$ in the forward pass and the spikes processed in each forward step. The number of steps scales linearly with the number of spikes processed in the forward pass, and it is important to understand that due to asynchronous processing the number of spikes processed until the evaluation of the loss function may be (well) less than the total number of spikes emitted throughout the forward pass. Since for every forward step, the computations are repeated, the time complexity scales linearly with the number of forward steps, $O(N_{t})$. The same applies to the space complexity.
During the backward pass, the spikes in $s_{i}$ that are not selected for processing can skip the computation $f$ for that step:
@@
Skip connections have been explored in deep ANNs and identified as a contributor to the stability of the training process Orhan & Pitkow (2018). This may apply to the skip in unlayered backpropagation, as well. The extent to which this is the case remains unexplored in this work.
Note that under this generalization, layered backpropagation corresponds to $F$-group the size of a layer, populated with a static round-Robin execution schedule following neuron index order within a layer, and re-reset across consecutive layers).
Finally, to leverage some visual insight into what makes unlayered backpropagation different from canonical (layered) backpropagation as a consequence of the ramifications we have proposed, figure 3 illustrates the forward and backward paths in the two cases (the forward path in figure 3.c depicts essentially the control-flow of our simulator). The most notable difference is that because of the way activation flow is generated in a variable number of forward steps (which, unless $F$-group has equal size to the number of neurons in a layer, is often more than the number of layers in the network), the unrolled compute graph in the backward pass will be different at every timestep, which contrasts with the fixed lattice-like structure of the compute graph in canonical backpropagation. This results in (fundamentally) different gradient flow and learning a different model (more robust for asynchronous inference, as our experiments will show).
Figure 3: (a) The forward pass with layer synchronization, where layer $0\leq j<L$ receives spike input $s_{j}$, creating currents that are integrated into membrane potentials $u_{j}$, which are carried over to the next timestep, and potentially emitting new spikes to the next layer if exceeding the spike threshold. (b) The backward pass of a single timestep for layered backpropagation. The structure of the computational graph is fixed across timesteps for a network with constant depth. (c) The forward pass for unlayered backpropagation, where $s_{\text{selected}}^{i}$ represents a selection of spikes from anywhere in the network propagated during the $i$th forward step, and $U$ denotes the membrane potentials of all neurons in the network. The membrane potentials are carried over to the next timestep at the end of the forward pass. Spikes from the last layer are pulled out of the loop as output. (d) The backward pass for unlayered backpropagation of a single timestep that consisted of $n+1$ forwards steps. Unlike layered backpropagation, the structure of the computational graph varies across timesteps.
@@ 4.1 Experimental setup
We carried out our experiments on SNN models trained primarily in three common benchmarking datasets, each of them has a different structure: N-MNIST Orchard et al. (2015), SHD Cramer et al. (2020), and DVS gestures Amir et al. (2017). N-MNIST has purely spatial structure, SHD purely temporal, and DVS gestures combines both spatial and temporal (input framing in DVS gesture is done such that an entire gesture motion and contour is not revealed within a single frame). These three datasets were processed with fully connected network topologies, without explicit lateral connectivity. The input was provided sequentially across a number of timesteps, allowing the networks to operate recurrently, exploiting their statefulness (implicit recurrence) and neuron state decay, both of which are needed for temporal tasks. We also repeated some of the experiments with a fourth dataset, CIFAR-10 Krizhevsky (2012), in a more complex and deep convolutional VGG-style network. In this case, as the input consists of static images, it was fed to the network in a single timestep. Our experiments involved primarily models with rate coding without and with refractoriness forcing neurons down to few/single spikes per timestep. We did not experiment with population coding. More details on the datasets and network topologies can be found in sections A.6.1,A.10. Table 1 summarizes the parameterization of the experiments and the reported results. The network architecture and hyperparameters are given in table 6. State-of-the art performance for these tasks can be achieved with reasonably shallow and wide models. We chose however to train narrow, but deeper network architectures so that the effects of absence of layer synchronization can be revealed in the comparison (because of this bias in network topology the accuracy results shown can vary from the top state of the art). In section A.7, we also provide an additional ablation with regard to how various training hyperparameters affect accuracy (forward group size, refractory dropout, and momentum noise during training). Results on error convergence are provided in section A.9.
@@
Stop condition Forward pass terminates: If all network activity drains ("On spiking done") or one forward step after the first spike is emitted by the output layer ("On output"). The default is "On spiking done" during training and "On output" during inference.
@@ 4.2 Network asynchrony increases neuron reactivity
One expects that more reactive neurons imply higher activation density. Interestingly, this is not necessarily the case for the models trained for asynchronous inference! In figure 4 (bottom row) we see that if we wait for the network to “drain” of spike activity during a forward pass the total number of spikes will indeed be higher. But if the forward pass terminates as soon as a decision is made, asynchronous models are consistently sparser. This is because asynchronous processing allows spike activity to freely flow through to the output and not be blocked at every layer for synchronization.
4.3 Unlayered backpropagation recovers accuracy and increases sparsity
Network asynchrony negatively affects the performance of the models trained with layered backpropagation in all three datasets (table 2). This is likely the “Achilles’ heel” of neuromorphic processing today. However, we observe that the accuracy loss is remediated when training takes into account asynchronous processing dynamics, which also significantly increases sparsity (by about 22x). We witnessed that the accuracy of models trained and executed asynchronously is consistently superior under the two scheduling policies we considered. This result conjectures that neuromorphic AI is competitive and more computationally efficient.
@@ 4.3 Unlayered backpropagation recovers accuracy and increases sparsity
Figure 5: Accuracy as function of forward steps after the first spike in the output layer. Given that $F=8$, each extra forward step processes another 8 spikes, assuming enough spikes are available. Dashed lines show the accuracy after all spike activity has been “drained” out of the network.
4.4 Network asynchrony and energy efficiency
[…]
N-MNIST SHD DVS gesture
Training Inference SOs Energy SOs Energy SOs Energy
Layered Layered 3.521 0.9155 50.82 13.21 158.8 41.29
Layered Async RS 3.225 0.8386 49.12 12.77 150.9 39.22
Unlayered RS Async RS 1.328 0.3454 18.02 4.684 85.92 22.34
Unlayered MS Async MS 1.304 0.3389 21.99 5.717 91.2 23.71
4.5 Network asynchrony reduces latency
An equally important result concerns the inference latency reduction under asynchronous network processing. The models trained with unlayered backpropagation have significantly lower latency than those trained with layer synchronization. Assuming a unit latency for processing one spike, the comparatively worst case latency will be given under sequential processing, as a function of the number of spikes processed until an inference decision is made. Figure 6 shows a distribution of the inference latencies across the entire test-set. It should be clear by now that this is because “the important spikes” in these models quickly reach the output layer, uninhibited by layer synchronization.
@@ 4.5 Network asynchrony reduces latency
4.6 Unlayered backpropagation is (currently) resource-intensive
We also tried to confirm the results in a scaled up setup, namely with a deeper VGG-7 like network (details in A.10), trained on the CIFAR-10 dataset. Figure 7a confirms the problem when training with layered backpropagation and the running asynchronous inference. When we try to train asynchronous models with unlayered backpropagation and the simpler RS scheduling policy the memory cost however becomes prohibitive unless we substantially increase the forward group size $F$. Nevertheless, even with as large as $F=512$ (during training, the smallest possible with an NVIDIA RTX 5000 GPU) and tested with $F=64$, accuracy steeply recovers to 71%, while activation density drops to about $1/4$ (confirming our previous observations). We anticipate that with much smaller $F$ during training (higher degree of network asynchrony), the accuracy can be completely recovered.
Unfortunately, the computational cost of unlayered backpropagation, in the current framework (and implementation) is rather high, especially when training with a small $F$. Each time $F$ is halved, the time and memory requirements approximately double as discussed in sections 3.2.1 and A.8. While this computational cost imposes obvious scalability constraints in the current implementation of unlayered-backpropagation for training very deep models with complex structures, it does not burden however inference on-chip. But it creates a trade-off, since the costlier computationally the training becomes, the more robust asynchronous inference will be when the resulting model is deployed on-chip.
CIFAR-10
Training Inference Acc. Density
Layered Layered 0.805 0.0089
Layered Async RS 0.183 0.0111
Async RS Async RS 0.712 0.0026
Figure 7: (a) Impact of network asynchrony on a CIFAR-10 trained VGG-style model. A larger decrease in accuracy occurs with smaller forward group sizes accompanied by an increase in activations. (b) Accuracy and activation density results on CIFAR-10.
5 Discussion & Future outlook
We pinpoint a crux for neuromorphic AI processing in training SNN models conventionally with backpropagation and naively assuming they will execute consistently, and energy and latency efficiently in neuromorphic processors. The problem lies in neglecting the dynamics of asynchronous processing, and we found a way to factor them in the gradient training process. Resulting models not only recover the affected performance (and even exceed it) but also exhibit the energy and latency benefits touted by neuromorphic computing.
Asynchronous processing is mainly about neurons acting independently of other neurons (also across layer boundaries) based on locally emerging flow dynamics, which is feasible with rate-like codings. So long as these dynamics are unobstructed by synchronization barriers, there is no need for explicit time-tracking. Temporal coding schemes requiring explicit time tracking (timestamps) typically do so due to the presence of synchronization points in event processing. Moreover, order coding and TTFS schemes which may use timestamps to enforce strict per/across-layer ordering, can also do without them by working with queues. The overarching assumption is that exact timing is relevant mainly to the extent that it facilitates the relative ordering of events.
From this viewpoint, in our framework synaptic currents arriving at a neuron trigger instantaneous evaluation of the membrane threshold independently of other incoming currents (i.e. there is no integration time interval at the neuron level). This eliminates the need to keep an explicit global clock for spike times. The instantaneous membrane voltage of neurons then reflects the temporal dynamics of the incoming activations at the neuron level. The use of dynamic scheduling allows neuron order dynamics to be reflected in spike exchanges without synchronization delays at layer boundaries. In other words, these modeling ramifications reflect relative temporal (order) dynamics among neurons across layers, unfolding inside the entire network under the influence of (a) the input distribution, and (b) the intrinsic operation of the underlying computing substrate (e.g., AI accelerator).
Accounting for neuron execution scheduling strategies in-training is likely a missing link for between neural algorithms/models and neuromorphic hardware architecture co-optimization. This study merely scratched the surface of a research exploration in this direction, but it shows that unless we rethink our training methods to account for asynchronous dynamics and redesign neuromorphic accelerators to move away from layer synchronization primitives, SNNs and brain-inspired computing could fail to deliver both performance and efficiency.
It is worth noting that the discrepancy between the dynamics encoded in neurons’ membranes and the way these dynamics are propagated across the neural network in current SNNs, was also independently identified and published only recently in Yao et al. (2024; 2025). To address this perceived as "modeling flaw" in how SNNs work, they propose training ANN models and converting them to SNNs by translating quantized integer activations to spike bursts of equivalent cardinality; or directly training bespoke SNN models where neurons fire not single spikes but spike bursts proportional to their membrane potentials. While these approaches seem to at least preserve ANN performance, they are only suited for models of integrate-and-fire (IF) neurons, since membrane leakage in LIF neurons would become ineffective with bursts of this sort. More importantly, the spike rates increase dramatically with adverse effect on energy consumption (which is a function of the spike density). By contrast, in the herein presented exploration we identify the "flaw" in the per-layer synchronization mechanism (which is a baggage from ANNs, but not inherently native to SNNs) that blocks these dynamics from expressing freely throughout the network. We show that casting away layer synchronization and enabling dynamics driven, adaptive neuron-execution scheduling during training per-se, we can train models aware of various asynchronous hardware, that finish inference faster than hardware agnostic ones, and provide correct predictions by processing only a subset of the spikes (improving both on energy and latency efficiency). Either approach however is in infancy, deserving more research and engineering for addressing scalability and demonstrating practical testing and efficiency with modern large deep models (e.g. as proposed in Yao et al. (2025)) on neuromorphic accelerators that can accommodate them (e.g. Davies et al. (2018); Furber et al. (2014); Tang et al. (2023)).
This motivates our follow-up work on challenges along 2 parallel directions: (a) improving on the computational resource limitations of asynchronous training with unlayered backpropagation, and (b) bridging this work to more complex contemporary model structures and constructs (e.g. Deng et al. (2022); Yao et al. (2022); Guo et al. (2023b); Yu et al. (2022)). Regarding the former direction, recent advances with event-based back-propagationWunderlich & Pehle (2021a) and sparse gradient learning Lohoff et al. (2025), show promise for addressing these challenges.
Acknowledgments
The herein research was partially sponsored and funded by Imec Netherlands, and the EU’s Horizon Europe Research and Innovation programme (under Grant Agreement 101070679). Guangzhi Tang is partially funded by the Dutch Research Council’s programme AiNed XS Europe programme (under Grant agreement NGF.1609.243.044, https://doi.org/10.61686/MYMVX53467). Roel Koopman is funded by the Dutch Research Council (under Grant agreement KICH1.ST04.22.021).